Suppose a random variable is normally distributed. The probabilities for a85 and 142 are 10% and 65%, respectively. Find the mean and standard deviation to the nearest whole number.


1
Expert's answer
2021-12-06T17:09:00-0500

Mean (Expected Value):

"Mean (\u03bc) = \u03a3xp"

"Mean (μ) = (85\times0.10)+(142\times0.65)"

"Mean (μ) =8.50+92.30=100.80\approx101"


Standard Deviation (σ):

"\sigma=\sqrt{Σ(x-\mu)^{2} \rho}"


"\sigma=\sqrt{(85-101)^{2}\times0.1+(142-101)^2\times0.65}"

"\sigma=\sqrt{(256\times0.1)+(1681\times0.65)}"


"\sigma=\sqrt{25.6+168.1}"


"\sigma=\sqrt{193.7}"


"\sigma=13.9\approx14"